# Rethinking Individual Fairness in Deepfake Detection

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 33rd ACM International Conference on Multimedia; October 27 - October 31, 2025; Dublin, Ireland.isbn: 978-1-4503-XXXX-X/18/06ccs: Multimedia in the Generative AI Era Social Aspects of Generative AI
@@ 1. Introduction
Recently, numerous deepfake detectors based on DNNs have been developed, aiming to accurately capture spatial and frequency-based artifact characteristics and utilize these learned features to distinguish real from manipulated media (Sinitsa and Fried, 2024; Wang et al., 2023; Ma et al., 2023; Qian et al., 2020). Beyond pure detection, contemporary research increasingly emphasizes the enhancement of detectors with additional functionalities, including identifying the specific generative model used to create fake media (Guarnera et al., 2024; Sha et al., 2023), ensuring that detectors maintain generalization capabilities against unseen forgery methods (Yan et al., 2024; Zheng et al., 2024; Chen et al., 2024a, b), explaining detection outcomes, localizing manipulated regions within media (Xu et al., 2025; Hong et al., 2024a; Guo et al., 2023), and improving the reliability of detectors even when tested on media of lower quality than training data (Ju et al., 2023; Xu et al., 2023; Jer Wong and Ng, 2023; Hooda et al., 2024). Despite these advancements, fairness in deepfake detectors remains considerably less explored compared to other aspects. This oversight significantly limits detectors’ applicability, as attackers may exploit biases by targeting particular populations, thus circumventing detection systems.
Only a few recent studies have developed algorithmic approaches aimed at improving fairness in deepfake detection. For example, Ju et al.(Ju et al., 2024) introduced two fairness losses based on distributionally robust optimization and fairness measures. Both methods encourage similar loss values across demographic groups and demonstrate promising fairness outcomes under intra-domain testing, where the training and test data originate from the same domain. However, these approaches do not ensure fairness generalization in cross-domain evaluation, where test samples are generated from previously unseen forgery methods. To address this limitation, Lin et al.(Lin et al., 2024c) proposed a disentanglement learning framework specifically aimed at enhancing the generalization of fairness for deepfake detectors. Although both studies effectively improve fairness at the group level, the individual-level fairness is largely overlooked.
@@
    We propose the first general framework designed to flexibly and seamlessly integrate into a wide range of existing deepfake detection methods to effectively enhance their individual fairness as well as generalization capabilities.
@@ 2. Related Work
Deepfake Detection. Current deepfake detection methods can be categorized into three primary groups based on the features they employ. The first category hinges on identifying inconsistencies in the physical and physiological characteristics of deepfakes (Lin et al., 2024b). For example, inconsistent corneal specular highlights (Hu et al., 2021), the irregularity of pupil shapes (Guo et al., 2022a, b), eye blinking patterns (Li et al., 2018), eye color difference (Matern et al., 2019), facial landmark locations (Yang et al., 2019), etc. The second category concentrates on signal-level artifacts introduced during the synthesis process, especially those from the frequency domain (Qian et al., 2020). These methods encompass various techniques, such as examining disparities in the frequency spectrum (Khayatkhoei and Elgammal, 2022; Dzanic et al., 2020) and utilizing checkerboard artifacts introduced by the transposed convolutional operator (Frank et al., 2020; Zhang et al., 2019). However, the methods in the above two categories usually exhibit relatively low detection performance. Therefore, the largest portion of existing detection methods fall into the data-driven category, including (Guo et al., 2022c; Pu et al., 2022a; Yan et al., 2023a; Liu et al., 2024; Hong et al., 2024b; Choi et al., 2024). These methods leverage various types of DNNs trained on both authentic and deepfake videos to capture specific discernible artifacts. While these methods achieved strong detection utility, they overlooked detection fairness, resulting in large performance disparities among groups (e.g., gender).
@@
Fairness in Deepfake Detection. Many studies (Trinh and Liu, 2021; Hazirbas et al., 2021; Pu et al., 2022b; Xu et al., 2024; Lin et al., 2024d; Nadimpalli and Rattani, 2022) have identified fairness issues in deepfake detection. For instance, Trinh et al. (Trinh and Liu, 2021) found biases present in both datasets and detection models, resulting in significant disparities in error rates across demographic groups. Similar findings were reported by Hazirbas et al. (Hazirbas et al., 2021). Pu et al. (Pu et al., 2022b) evaluated the fairness of the MesoInception-4 model using the FF++ dataset, observing performance bias related to gender. Nadimpalli et al. (Nadimpalli and Rattani, 2022) also highlighted substantial biases and proposed a preprocessing method involving a gender-balanced dataset to mitigate gender-based performance disparities. However, this approach delivered only modest improvements and necessitated extensive data annotation efforts. To further support fairness-oriented algorithm development, (Xu et al., 2024) and (Lin et al., 2024d) performed comprehensive analyses of bias in deepfake detection, augmenting datasets with diverse demographic annotations. Recent works by Ju et al. (Ju et al., 2024) and Lin et al. (Lin et al., 2024c) have subsequently concentrated on enhancing group-level fairness through algorithmic interventions. However, these studies have not comprehensively examined fairness at the individual level with solutions, an aspect that remains largely unaddressed in recent deepfake detection research and constitutes the primary focus of our work.
Individual Fairness. To the best of our knowledge, individual fairness in deepfake detection remains unexplored. However, this challenge has been studied in broader machine learning contexts. The concept of individual fairness was first introduced by Dwork et al. (Dwork et al., 2012) under the principle that “similar individuals with respect to a particular task should be treated similarly.” The major challenge in individual fairness is the definition and estimation of an appropriate similarity metric, which remains a bottleneck in its broader application (Mukherjee et al., 2020). Several recent studies have extended individual fairness beyond its original definition. Kearns et al. (Sharifi-Malvajerdi et al., 2019) proposed average individual fairness, which balances statistical and individual fairness by ensuring fairness across both individuals and classification tasks. Mukherjee et al. (Mukherjee et al., 2020) presented two data-driven approaches to learn fair metrics, making individual fairness more practical to broader machine learning tasks. (Fleisher, 2021) argued that similarity-based fairness can still encode human biases, leading to systematic disparities if the metric itself is flawed. Furthermore, fairness constraints must be flexible enough to accommodate contextual factors, such as decision-dependent fairness trade-offs in high-stakes settings. Given these findings, applying individual fairness in deepfake detection requires rethinking similarity metrics. Unlike traditional image classification, where samples of the same class exhibit higher similarity based on observed semantics, deepfakes often exhibit greater similarity across different classes, as demonstrated in our following Motivation section. Consequently, the conventional principle of individual fairness cannot be directly applicable, requiring us to rethink individual fairness in deepfake detection.
@@ 4. The Naive Approach
Given a training dataset $\mathcal{S}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ with size $n$, where $X\in\mathbb{R}^{d}$ is the image with dimension $d$, $Y\in\{0,1\}$ is the corresponding label (0 for fake, 1 for real). A straightforward method to enhance individual fairness in deepfake detection is to incorporate a fairness regularization term into a baseline classifier. Specifically, let us consider a deep learning model with a feature extractor $E(\cdot)$ and a classifier head $h(\cdot)$. The conventional training for a deepfake detector is to minimize the following learning objective: $\mathcal{L}_{CE}=\frac{1}{n}\sum_{i=1}^{n}C(h(E(X_{i})),Y_{i})$, where $C(\cdot,\cdot)$ represents the Cross-Entropy (CE) loss. To enforce individual fairness, a naive approach widely used in existing works (Ju et al., 2024; Lin et al., 2024c) is to introduce an additional fairness regularization term, $\mathcal{L}^{naive}_{ind}$, which penalizes discrepancies in model outputs for similar input images. The resulting objective function can be represented as:
@@ 5.1. Anchor Learning
First, we address the issue arising from the $\mathcal{L}_{CE}$ loss and the term $\spadesuit$. While existing approaches (Lin et al., 2024c; Ren et al., 2025; Yan et al., 2023a, 2024), such as disentanglement learning and data augmentation, have been proposed to enhance generalization in deepfake detection, these methods typically result in increased model complexity or larger training datasets. Consequently, they may not be applicable to integrate into advanced detection methods or restricted training environments. This motivates us to develop a more efficient and broadly generalizable approach.
@@ 5.2. Semantic-Agnostic Individually Fair Learning
Breaking Semantic Correlations. Since the high similarity in semantic features across different classes in deepfake datasets (as shown in Fig. 1 (a) and (b)) conflicts with the definition of individual fairness loss (see Eq. (1)), training a detector with the above naive approach fails to mitigate bias. This bias further arises as models overfit identity-related attributes rather than focus on the manipulation artifacts. To address this challenge, we employ a patch shuffle strategy. First, given an input image $X$, we divide it into multiple non-overlapping patches of size $P\times P$:
$\{p_{o}\}=\text{Partition}(X,P),\quad o=1,2,\dots,K\times K$
where $K\times K$ represents the total number of patches obtained after segmentation. To disrupt global semantic dependencies while preserving local details, we randomly shuffle these patches into a new sequence $\{p_{m}\},(m=1,2,\dots,K\times K)$. Finally, we reconstruct the image by rearranging the shuffled patches back to their original dimensions, forming $X^{\prime}$, which serves as input for subsequent processing. This transformation ensures that the model focuses on fine-grained manipulation artifacts rather than identity-related semantic structures.
Exposing Artifacts. Generative models typically introduce subtle but systematic artifacts during the synthesis process. These artifacts often remain concealed under high-level semantic details, making them difficult to discern when analyzing raw images directly. To isolate and amplify these telltale signs, we adopt a residual-based strategy (Corvi et al., 2023a). Specifically, after applying patch shuffling to an input image $X$, we generate a denoised version $D(X^{\prime})$ with a filter from (Zhang et al., 2017) to capture the smooth, semantic content. We then compute the residual $\overline{X}^{\prime}$ by subtracting the denoised result from $X^{\prime}$,
$\overline{X}^{\prime}=X^{\prime}-D(X^{\prime}).$
Crucially, we compute the residual after patch shuffling, thereby exposing the remaining synthesized artifacts once semantic connections are disrupted.
@@ 5.3. Optimization
To improve generalization and avoid suboptimal solutions common in overparameterized DNNs, we apply the sharpness-aware minimization (SAM) technique (Foret et al., 2020) to flatten the loss landscape, which has been proven to efficiently improve generalization (Lin et al., 2024c; Ren et al., 2025; Lin et al., 2024a). Specifically, denoting the model weights of the whole framework as $\theta$, this involves finding an optimal $\epsilon^{*}$ to perturb $\theta$ in a way that maximizes the loss, expressed as:
@@
Here, $\gamma$ is a hyperparameter that controls the perturbation magnitude, and $\nabla_{\theta}\mathcal{L}$ is the gradient of $\mathcal{L}$ with respect to $\theta$. The approximation term is derived using a first-order Taylor expansion, assuming $\epsilon$ is small. The final equation is obtained by solving a dual norm problem. Thus, the model weights are updated by solving:
(6) $\min_{\theta}\mathcal{L}\textbf{(}\theta+\epsilon^{*}\textbf{)}.$
The underlying idea is that perturbing the model in the direction of the gradient norm increases the loss value, thereby improving generalization. We optimize Eq. (6) using stochastic gradient descent, as detailed in Algorithm 1 of the Appendix.
@@ 6.1. Experimental Settings
Datasets. To comprehensively assess the proposed method, we utilize five widely-used public datasets: FaceForensics++ (FF++) (Rossler et al., 2019), Deepfake Detection (DFD) (Google and Jigsaw, 2019), Deepfake Detection Challenge (DFDC) (Brian et al., 2020), Celeb-DF (Li et al., 2020), and the latest million-scale dataset, AI-Face (Lin et al., 2024d). AI-Face notably includes face images generated by advanced diffusion models, which are absent from the other four datasets, enabling the evaluation of our method against more sophisticated generative techniques. For intra-domain evaluation, we train models on each dataset’s training set and evaluate them on its corresponding test set. For cross-domain evaluation, following common practice in previous works (Ju et al., 2024; Lin et al., 2024c; Yan et al., 2023a), we train our model on FF++ and test it on the remaining datasets. Since we compare against existing fairness-enhanced methods (Ju et al., 2024; Lin et al., 2024c), we adopt their data processing and annotation settings to ensure fair comparisons. Details of each dataset are provided in Appendix 9.1. We conduct experiments using three popular backbones: Xception (Chollet, 2017), ResNet-50 (He et al., 2016), and EfficientNet-B3 (Tan and Le, 2019). Unless explicitly mentioned, all methods are employed on Xception (Chollet, 2017) backbone.
Evaluation Metrics. For detection utility, we use the Area Under the Curve (AUC), consistent with prior works (Ju et al., 2024; Lin et al., 2024c; Yan et al., 2023a). To evaluate individual fairness, we treat $\mathcal{L}_{\text{ind}}^{naive}$ and $\mathcal{L}_{\text{ind}}^{*}$ as the traditional and proposed individual fairness metrics, respectively.
Baseline Methods. To clearly demonstrate our method’s effectiveness in improving individual fairness in deepfake detection, we first define two fundamental baselines: (1) Ori, a backbone (e.g., Xception (Chollet, 2017)) trained solely with standard CE loss, and (2) Naive, the Ori detector trained using CE loss with traditional individual fairness regularization $\mathcal{L}_{ind}^{naive}$. Additionally, to illustrate our method’s plug-and-play flexibility, we integrate it into various state-of-the-art detectors across multiple categories, including spatial-based (UCF (Yan et al., 2023a), CORE (Ni et al., 2022), RECCE (Cao et al., 2022)), frequency-based (F3Net (Qian et al., 2020), SPSL (Liu et al., 2021), SRM (Luo et al., 2021)), and fairness-enhanced detectors (DAW-FDD (Ju et al., 2024), DAG-FDD (Ju et al., 2024), PG-FDD (Lin et al., 2024c)).

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755244isbn: 979-8-4007-2035-2/2025/10ccs: Security and privacy Human and societal aspects of security and privacyccs: Computing methodologies Computer vision
@@ 1. Introduction
Recently, numerous deepfake detectors based on DNNs have been developed, aiming to accurately capture spatial and frequency-based artifact characteristics and utilize these learned features to distinguish real from manipulated media (Sinitsa and Fried, 2024; Wang et al., 2023a; Ma et al., 2023; Qian et al., 2020; Krubha et al., 2025; Wang et al., 2024; Lin et al., 2024a; Zhang et al., 2024; Chen et al., 2024a; Lin et al., 2024c; Fan et al., 2023; Zhang et al., 2023; Pu et al., 2022a; Guo et al., 2022b, c; Wang et al., 2023b). Beyond pure detection, contemporary research increasingly emphasizes the enhancement of detectors with additional functionalities, including identifying the specific generative model used to create fake media (Guarnera et al., 2024; Sha et al., 2023), ensuring that detectors maintain generalization capabilities against unseen forgery methods (Yan et al., 2024; Zheng et al., 2024; Chen et al., 2024b, c; Ren et al., 2024; Yang et al., 2023), explaining detection outcomes, localizing manipulated regions within media (Xu et al., 2025; Hong et al., 2024a; Guo et al., 2023), and improving the reliability of detectors even when tested on media of lower quality than training data (Ju et al., 2023; Xu et al., 2023; Jer Wong and Ng, 2023; Hooda et al., 2024). Despite these advancements, fairness in deepfake detectors remains considerably less explored compared to other aspects. This oversight significantly limits detectors’ applicability, as attackers may exploit biases by targeting particular populations, thus circumventing detection systems.
Only a few studies (Wu et al., 2025; Ju et al., 2024; Lin et al., 2024d) have developed algorithmic approaches aimed at improving fairness in deepfake detection. For example, (Ju et al., 2024) introduced two fairness losses based on distributionally robust optimization and fairness measures. Both methods encourage similar loss values across demographic groups and demonstrate promising fairness outcomes under intra-domain testing, where the training and test data originate from the same domain. However, these approaches do not ensure fairness generalization in cross-domain evaluation, where test samples are generated from previously unseen forgery methods. To address this limitation, Lin et al.(Lin et al., 2024d) proposed a disentanglement learning framework specifically aimed at enhancing the generalization of fairness for deepfake detectors. Although both studies effectively improve fairness at the group level, the individual-level fairness is largely overlooked.
@@
    We propose the first general framework designed to seamlessly integrate into a wide range of existing deepfake detection methods, effectively enhancing their individual fairness as well as generalization capabilities.
@@ 2. Related Work
Individual Fairness. To the best of our knowledge, individual fairness in deepfake detection remains unexplored. However, this challenge has been studied in broader machine learning contexts. The concept of individual fairness was first introduced by Dwork et al. (Dwork et al., 2012) under the principle that “similar individuals with respect to a particular task should be treated similarly.” The major challenge in individual fairness is the definition and estimation of an appropriate similarity metric, which remains a bottleneck in its broader application (Mukherjee et al., 2020). Several recent studies have extended individual fairness beyond its original definition. Kearns et al. (Sharifi-Malvajerdi et al., 2019) proposed average individual fairness, which balances statistical and individual fairness by ensuring fairness across both individuals and classification tasks. Mukherjee et al. (Mukherjee et al., 2020) presented two data-driven approaches to learn fair metrics, making individual fairness more practical to broader machine learning tasks. However, (Fleisher, 2021) argued that similarity-based fairness can still encode human biases, leading to systematic disparities if the metric itself is flawed. Furthermore, unlike traditional image classification, where samples of the same class exhibit higher similarity based on observed semantics, deepfakes often exhibit greater similarity across different classes, as demonstrated in our following Motivation section. Consequently, the conventional principle of individual fairness is not directly applicable, requiring us to rethink individual fairness in deepfake detection.
@@ 4. The Naive Approach
Given a training dataset $\mathcal{S}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ with size $n$, where $X\in\mathbb{R}^{d}$ is the image with dimension $d$, $Y\in\{0,1\}$ is the corresponding label (0 for fake, 1 for real). A straightforward method to enhance individual fairness in deepfake detection is to incorporate a fairness regularization term into a baseline classifier. Specifically, let us consider a deep learning model with a feature extractor $E(\cdot)$ and a classifier head $h(\cdot)$. The conventional training for a deepfake detector is to minimize the following learning objective: $\mathcal{L}_{CE}=\frac{1}{n}\sum_{i=1}^{n}C(h(E(X_{i})),Y_{i})$, where $C(\cdot,\cdot)$ represents the Cross-Entropy (CE) loss. To enforce individual fairness, a naive approach widely used in existing works (Ju et al., 2024; Lin et al., 2024d) is to introduce an additional fairness regularization term, $\mathcal{L}^{naive}_{ind}$, which penalizes discrepancies in model outputs for similar input images. The resulting objective function can be represented as:
@@ 5.1. Anchor Learning
First, we address the issue arising from the $\mathcal{L}_{CE}$ loss and the term $\spadesuit$. While existing approaches (Lin et al., 2024d; Ren et al., 2024; Yan et al., 2023a, 2024), such as disentanglement learning and data augmentation, have been proposed to enhance generalization in deepfake detection, these methods typically result in increased model complexity or larger training datasets. Consequently, they may not be applicable to integrate into advanced detection methods or restricted training environments. This motivates us to develop a more efficient and broadly generalizable approach.
@@ 5.2. Semantic-Agnostic Individually Fair Learning
Breaking Semantic Correlations. Since the high similarity in semantic features across different classes in deepfake datasets (as shown in Fig. 1 (a) and (b)) conflicts with the definition of individual fairness loss (see Eq. (1)), training a detector with the above naive approach fails to mitigate bias. This bias further arises as models overfit identity-related attributes rather than focus on the manipulation artifacts. To address this challenge, we employ a patch shuffle strategy. First, given an input image $X$, we divide it into multiple non-overlapping patches of size $P\times P$: $\{p_{o}\}=\text{Partition}(X,P),\quad o=1,2,\dots,K\times K$ where $K\times K$ represents the total number of patches obtained after segmentation. To disrupt global semantic dependencies while preserving local details, we randomly shuffle these patches into a new sequence $\{p_{m}\},(m=1,2,\dots,K\times K)$. Finally, we reconstruct the image by rearranging the shuffled patches back to their original dimensions, forming $X^{\prime}$, which serves as input for subsequent processing. This transformation ensures that the model focuses on fine-grained manipulation artifacts rather than identity-related semantic structures.
Exposing Artifacts. Generative models typically introduce subtle but systematic artifacts during the synthesis process. These artifacts often remain concealed under high-level semantic details, making them difficult to discern when analyzing raw images directly. To isolate and amplify these telltale signs, we adopt a residual-based strategy (Corvi et al., 2023a). Specifically, after applying patch shuffling to an input image $X$, we generate a denoised version $D(X^{\prime})$ with a filter from (Zhang et al., 2017) to capture the smooth, semantic content. We then compute the residual $\overline{X}^{\prime}$ by subtracting the denoised result from $X^{\prime}$: $\overline{X}^{\prime}=X^{\prime}-D(X^{\prime})$. Crucially, we compute the residual after patch shuffling, thereby exposing the remaining synthesized artifacts once semantic connections are disrupted.
@@ 5.3. Optimization
To improve generalization and avoid suboptimal solutions common in overparameterized DNNs, we apply the sharpness-aware minimization (SAM) technique (Foret et al., 2020) to flatten the loss landscape, which has been proven to efficiently improve generalization (Lin et al., 2024d; Ren et al., 2024; Lin et al., 2024b). Specifically, denoting the model weights of the whole framework as $\theta$, this involves finding an optimal $\epsilon^{*}$ to perturb $\theta$ in a way that maximizes the loss, expressed as:
@@
Here, $\gamma$ is a hyperparameter that controls the perturbation magnitude, and $\nabla_{\theta}\mathcal{L}$ is the gradient of $\mathcal{L}$ with respect to $\theta$. The approximation term is derived using a first-order Taylor expansion, assuming $\epsilon$ is small. The final equation is obtained by solving a dual norm problem. Thus, the model weights are updated by solving: $\min_{\theta}\mathcal{L}\textbf{(}\theta+\epsilon^{*}\textbf{)}\end{aligned}$. The underlying idea is that perturbing the model in the direction of the gradient norm increases the loss value, thereby improving generalization. We optimize the SAM objective using stochastic gradient descent, as detailed in Algorithm 1 of the Appendix.
@@ 6.1. Experimental Settings
Datasets. To comprehensively assess the proposed method, we utilize five widely-used public datasets: FaceForensics++ (FF++) (Rossler et al., 2019), Deepfake Detection (DFD) (Google and Jigsaw, 2019), Deepfake Detection Challenge (DFDC) (Brian et al., 2020), Celeb-DF (Li et al., 2020), and the latest million-scale dataset, AI-Face (Lin et al., 2024e). AI-Face notably includes face images generated by advanced diffusion models, which are absent from the other four datasets, enabling the evaluation of our method against more sophisticated generative techniques. For intra-domain evaluation, we train models on each dataset’s training set and evaluate them on its corresponding test set. For cross-domain evaluation, following common practice in previous works (Ju et al., 2024; Lin et al., 2024d; Yan et al., 2023a), we train our model on FF++ and test it on the remaining datasets. Since we compare against existing fairness-enhanced methods (Ju et al., 2024; Lin et al., 2024d), we adopt their data processing and annotation settings to ensure fair comparisons. Details of each dataset are provided in Appendix 10.1. We conduct experiments using three popular backbones: Xception (Chollet, 2017), ResNet-50 (He et al., 2016), and EfficientNet-B3 (Tan and Le, 2019). Unless explicitly mentioned, all methods are employed on Xception (Chollet, 2017) backbone.
Evaluation Metrics. For detection utility, we use the Area Under the Curve (AUC), consistent with prior works (Ju et al., 2024; Lin et al., 2024d; Yan et al., 2023a). To evaluate individual fairness, we treat $\mathcal{L}_{\text{ind}}^{naive}$ and $\mathcal{L}_{\text{ind}}^{*}$ as the traditional and proposed individual fairness metrics, respectively.
Baseline Methods. To clearly demonstrate our method’s effectiveness in improving individual fairness in deepfake detection, we first define two fundamental baselines: (1) Ori, a backbone (e.g., Xception (Chollet, 2017)) trained solely with standard CE loss, and (2) Naive, the Ori detector trained using CE loss with traditional individual fairness regularization $\mathcal{L}_{ind}^{naive}$. Additionally, to illustrate our method’s plug-and-play flexibility, we integrate it into various state-of-the-art detectors across multiple categories, including spatial-based (UCF (Yan et al., 2023a), CORE (Ni et al., 2022), RECCE (Cao et al., 2022)), frequency-based (F3Net (Qian et al., 2020), SPSL (Liu et al., 2021), SRM (Luo et al., 2021)), and fairness-enhanced detectors (DAW-FDD (Ju et al., 2024), DAG-FDD (Ju et al., 2024), PG-FDD (Lin et al., 2024d)).